OpenStudy (anonymous):

Hey can anyone help??? how do you find the center and the radius of this equation, (x-1)^2+y^2=24

terenzreignz (terenzreignz):

General form of the equation of a circle is: \[(x-h)^{2}+(y-k)^{2}=r^{2}\] Where (h,k) is the center and the radius is r Now, what values do you have for h, k, and r? :)

OpenStudy (ghazi):

center= (1,0) and radius =\[\sqrt{24}\]

OpenStudy (anonymous):

how am i suppose to put it into standard form if i only have y^2 and no other y values??

OpenStudy (ghazi):

(x-1)^2+(y-0)^2=24

OpenStudy (anonymous):

oh so you sub 0 in if you dont have the other value you need??

OpenStudy (ghazi):

yes

OpenStudy (anonymous):

ok that makes much more sense.